# Uniqueness of sign-changing solutions to Trudinger’s equation

### v1

@@ Abstract.
We establish uniqueness for sign-changing solutions to Trudinger’s parabolic equation with time dependent $C^{2}$ Dirichlet boundary data.
@@ 1. Introduction
Especially, under natural assumptions the uniqueness of sign-changing solutions is a challenging open problem for Trudinger’s equation. For solutions $u$ with $\inf\{u\}>0$ this is settled, see [LL22] for references.
We shall focus on the uniqueness of sign-changing weak solutions belonging to the Sobolev space $L^{p}(0,T;W^{1,p}(\Omega))$. Our main result, Theorem 4.1, establishes uniqueness for sign-changing solutions with $C^{2}$ boundary data. For solutions with time-independent Dirichlet boundary values on the lateral boundary $\partial\Omega\times[0,T]$ F. Otto proved uniqueness in [Ott96]: for example weak solutions in $L^{p}(0,T;W^{1,p}_{0}(\Omega))$ are unique. In fact, these solutions turn out to have a time derivative $\partial_{t}(|u|^{p-2}u)$. Less regular solutions in $L^{p}(0,T;W^{1,p}_{0}(\Omega))$ cannot exist! We extend Otto’s result to time dependent boundary data, yet of class $C^{2}$. The idea of our proof is, first following Otto’s approach in [Ott96], to use Kružkov’s [Kru70] doubling of time variables to establish the comparison principle, but now extended to time dependent boundary values. The argument requires some space between the boundary values. Thus an extra device is needed: we finish the proof by considering the auxiliary solutions
@@
We can conclude the proof by using the $C^{2}$ assumption as in [LPS25a] to prove that both auxiliary solutions converge to the same solution as $\varepsilon\to 0$.
It has not escaped our attention that the $C^{2}$ assumption can be somewhat relaxed. Attempts to prove Theorem 3.1 for $\varepsilon=0$ have hitherto not succeeded.
@@ 2. Preliminaries
Throughout the paper, $\Omega\subseteq{\mathbb{R}}^{n}$ is a bounded Lipschitz domain, and we consider the fixed parameters $1<p<\infty$ and $0<T<\infty$. We use the standard notation $\Omega_{T}:=\Omega\times(0,T)$,
$\partial_{p}\Omega_{T}:=(\overline{\Omega}\times\{0\})\cup(\partial\Omega\times[0,T]),$
and denote the parabolic Sobolev space by $L^{p}(0,T;W^{1,p}(\Omega))$. Recall that this function space consists of measurable functions $u:\Omega_{T}\to{\mathbb{R}}$ such that the function $x\mapsto u(x,t)$ belongs to $W^{1,p}(\Omega)$ for almost every $t\in(0,T)$, and $u$ has the finite norm
$\biggl(\int\limits_{0}^{T}\!\!\int\limits_{\Omega}\left|u\right|^{p}+\left|\nabla u\right|^{p}\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t\biggr)^{\frac{1}{p}}<\infty.$
Here $\nabla u$ denotes the spatial gradient. The function space $L^{p}(0,T;W^{1,p}_{0}(\Omega))$ is defined analogously. We use the function $b$ defined by
$b(z):=\left|z\right|^{p-2}z,\qquad b(0)=0$
for real values. We write $z^{+}=\max\{z,0\}$.
@@
$u\in L^{p}(0,T,W^{1,p}(\Omega))$. We say that $u$ is a weak solution with the (Dirichlet) boundary values $\psi$ if the following conditions hold.
@@ 3. Comparison principle
It will be important not to assume any existence of time derivatives. Indeed, under such an extra assumption the result is well-known. For the convenience of the reader, we present a standard proof following the notation in [LPS25a]: Consider two solutions $u_{1},u_{2}$ that have time derivatives $\partial_{t}(|u_{j}|^{p-2}u_{j})$ at least in $L^{1}(\Omega_{T})$ and the same boundary values. Let $H_{\delta}(s)$ denote the approximation
@@
of the Heaviside function. We use the test function $\zeta=H_{\delta}\bigl(u_{2}-u_{1}\bigr)$ in the equation
@@
Next we state the comparison principle with boundary values
$\psi\in C(\overline{\Omega_{T}})\cap L^{p}(0,T;W^{1,p}(\Omega))$. Observe that now no existence of time derivatives for solutions is assumed.
@@
Let $u_{1}$ and $u_{2}$ be two weak solutions with boundary values $\psi_{1},\psi_{2}$ as in Definition 2.1. Assume that there is a constant $\varepsilon>0$ such that
(3.3) $\psi_{1}(x,t)\leq\psi_{2}(x,t)-\varepsilon\quad\text{ for all }\quad(x,t)\in\partial_{p}\Omega.$
@@
$q_{\sigma,+}(z_{1},w)-q_{\sigma,+}(z_{2},w)$ $=\,\,\int\limits_{z_{1}}^{z_{2}}\eta_{\sigma,+}^{\prime}(s-w)\,b^{\prime}(s)\,\text{d}s$
$\geq\,\,\eta_{\sigma,+}^{\prime}(z_{2}-w)\int\limits_{z_{1}}^{z_{2}}\,b^{\prime}(s)\,\text{d}s$ $=\,\,\eta_{\sigma,+}^{\prime}(z_{2}-w)\,(b(z_{1})-b(z_{2})).$
@@
We sometimes omit the spatial variable $x$ and write $u(t)=u(x,t)$ and $\nabla u(t)=\nabla u(x,t).$ Finally, let $\phi\in C_{0}^{\infty}({\mathbb{R}})$ be a function satisfying
@@
For a parameter $\lambda>0$ and a non-negative function $\gamma\in C_{0}^{\infty}(0,T)$, we define the function $\gamma_{\lambda}\in C_{0}^{\infty}({\mathbb{R}}^{2})$ by
@@
$|\psi_{1}(x,t)-\psi_{2}(x,s)|<\varepsilon/2\quad\text{when}\quad\left|t-s\right|<\delta$
@@
Recall that $\eta_{\sigma,+}(z)=0$ when $z\leq 0$. We next observe that
@@
$+$ $\gamma_{\lambda}(t,s)\eta_{\sigma,+}^{\prime\prime}(u_{1}(t)\!-\!u_{2}(s))\langle\left|\nabla u_{1}(t)\right|^{p-2}\nabla u_{1}(t),\nabla u_{1}(t)-\nabla u_{2}(t)\rangle\Big\}\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t.$
@@
As for the proof of (3.7), note that $\eta_{\sigma,-}(z)=0$ when $z\geq 0$. Hence, by a similar argument as in the beginning of the proof, we have that
$g(x,s):=\eta_{\sigma,-}(u_{2}(s)-u_{1}(t))\,\gamma_{\lambda}(t,s)$
@@
$0\,\geq\int\limits_{0}^{T}\!\!\!\int\limits_{0}^{T}\!\!\!\int\limits_{\Omega}\!$ $\Bigl\{-q_{\sigma,+}(u_{1}(t),u_{2}(s))\partial_{t}\gamma_{\lambda}(t,s)-q_{\sigma,-}(u_{2}(t),u_{1}(s))\partial_{s}\gamma_{\lambda}(t,s)\Bigr$.
@@
$\int\limits_{0}^{T}\!\!\!$ $\int\limits_{0}^{T}\!\!\!\int\limits_{\Omega}\!\left(q_{\sigma,+}(u_{1}(t),u_{2}(s))\partial_{t}\gamma_{\lambda}(t,s)+q_{\sigma,-}(u_{2}(t),u_{1}(s))\partial_{s}\gamma_{\lambda}(t,s)\right)\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t\hskip 0.50003pt\textup{d}s\,$
@@
We now prove (3.8) by letting $\sigma\to 0+$. To this end, we verify a suitable convergence. First, note that when $\sigma\to 0+$
@@
$q_{\sigma,-}(z,w)\to(b(w)-b(z))^{+}\quad\text{ and }\quad\left|q_{\sigma,-}(z,w)\right|\leq(b(z)-b(w))^{+}.$
Hence, by letting $\sigma\to 0+$, inequality (3.8) follows from Lebesgue’s dominated convergence theorem. ∎
Next, we let $\lambda\to 0+$ in (3.8) to prove the following estimate.
@@
By a standard procedure for approximating Dirac’s delta with mollifiers, letting $\lambda\to 0+$ we arrive at
@@
See Lemma 2.9 on page 76 in [HH15] for details. This completes the proof. ∎
@@
Since
$\int\limits_{\Omega}\left(b(u_{1}(t_{2}))-b(u_{2}(t_{2}))\right)^{+}\!\hskip 0.50003pt\textup{d}x\leq$ $\int\limits_{\Omega}\left(b(u_{1}(t_{1}))-b(\psi_{1}(t_{1})\right)^{+}\!\hskip 0.50003pt\textup{d}x$
$+\int\limits_{\Omega}\left(b(\psi_{1}(t_{1}))-b(\psi_{2}(t_{1})\right)^{+}\!\hskip 0.50003pt\textup{d}x$
$+\int\limits_{\Omega}\left(b(u_{2}(t_{1}))-b(\psi_{2}(t_{1}))\right)^{+}\!\hskip 0.50003pt\textup{d}x.$
@@ 4. Proof of the uniqueness
We now note that there exists one and only one such $t$-regular solution, see [LPS25a, Pages 4-6] for details. They are unique at least within their own regularity class. The next theorem guarantees that there are no other weak solutions
@@
A solution with boundary values $\psi\in C^{2}(\overline{\Omega_{T}})$ is unique.
@@
Let $u$ be an arbitrary solution with the given boundary values $\psi$. According to [LPS25a, Theorem 5], there always exists a $t$-regular solution with the boundary values $\psi\in C^{2}(\overline{\Omega_{T}})$. Moreover, it is well-known that such a $t$-regular solution is unique, see for example [LPS25a, Corollary 3], although the existence of less regular solutions is not yet excluded. Given a constant $\varepsilon>0$, there exists unique $t$-regular solutions
@@
The proof has an interesting consequence. Since we obtained $u=u_{0}$, we can read off the following result from the estimate in [LPS25a, Theorem 5].
@@
$\partial_{t}(|u|^{\frac{p-2}{2}}u)\,\in L^{2}(\Omega_{T}),\qquad p\geq 2$
$\partial_{t}(|u|^{p-2}u)\,\in\,L^{p}(\Omega_{T}),\qquad 1<p<2.$

### v2

@@ Abstract.
We establish uniqueness of sign-changing solutions to Trudinger’s parabolic equation with time dependent $C^{2}$ Cauchy-Dirichlet boundary data.
@@ 1. Introduction
Especially, under natural assumptions the uniqueness of sign-changing solutions is a challenging open problem for Trudinger’s equation. For solutions $u$ with $\inf\{u\}\geq 0$ this is settled, see [LL22, LPS25a] for references.
We shall focus on the uniqueness of sign-changing weak solutions with time dependent boundary data. Our main result, Theorem 4.1, establishes uniqueness for sign-changing solutions with $C^{2}$ boundary data. It has not escaped our attention that the $C^{2}$ assumption can be somewhat relaxed. Proposition 3.6 covers the special case of monotone lateral boundary values without additional regularity. For solutions with time-independent Dirichlet boundary values on the lateral boundary $\partial\Omega\times[0,T]$  F. Otto proved uniqueness in [Ott96]: for example weak solutions in $L^{p}(0,T;W^{1,p}_{0}(\Omega))$ are unique. In fact, these solutions turn out to have a time derivative $\partial_{t}(|u|^{p-2}u)$. Less regular solutions in $L^{p}(0,T;W^{1,p}_{0}(\Omega))$ cannot exist!
The idea of our proof is, first following Otto’s approach in [Ott96], to use Kružkov’s [Kru70] doubling of time variables to establish the comparison principle, but now extended to time dependent boundary values. The argument requires some space between the boundary values. Thus an extra device is needed: we finish the proof by considering the auxiliary solutions
@@
We can conclude the proof by using the $C^{2}$ assumption as in [LPS25a] to prove that both auxiliary solutions converge to the same solution as $\varepsilon\to 0$. The limit procedure is based on an estimate which has been worked out only for $p>2$. However, Sections 2 and 3 are valid in the full range $1<p<\infty$.
@@ 2. Preliminaries
Let $\Omega\subseteq{\mathbb{R}}^{n}$ be a bounded Lipschitz domain. We consider the equation in the domain $\Omega_{T}=\Omega\times(0,T)$, when $1<p<\infty$. (In Section 4 we must take $p\geq 2$.) We use the standard notation
$\partial_{p}\Omega_{T}:=(\overline{\Omega}\times\{0\})\cup(\partial\Omega\times[0,T]).$
for the parabolic boundary and denote the parabolic Sobolev space by $L^{p}(0,T;W^{1,p}(\Omega))$. Recall that this function space consists of measurable functions $u:\Omega_{T}\to{\mathbb{R}}$ such that the function $x\mapsto u(x,t)$ belongs to $W^{1,p}(\Omega)$ for almost every $t\in(0,T)$, and $u$ has the finite norm
$\biggl(\int\limits_{0}^{T}\!\!\int\limits_{\Omega}\bigl(\left|u\right|^{p}+\left|\nabla u\right|^{p}\bigr)\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t\biggr)^{\frac{1}{p}}<\infty.$
Here $\nabla u$ denotes the spatial gradient. The function space $L^{p}(0,T;W^{1,p}_{0}(\Omega))$ is defined analogously. Define
$b(z):=\left|z\right|^{p-2}z,\qquad b(0)=0.$
We write $z^{+}=\max\{z,0\}$.
@@
$u\in L^{p}(0,T;W^{1,p}(\Omega))$. We say that $u$ is a weak solution with the (Dirichlet) boundary values $\psi$ if the following conditions hold.
@@
Remark. It is customary to treat the (Cauchy) initial values and the (Dirichlet) lateral boundary values separately, as in [AL83], [Ott96], [BDL21] for instance. So we require
$\begin{cases}u(x,0)=\psi_{0}(x)\quad\text{when}\quad x\in\Omega\\ u(x,t)=\psi(x,t)\quad\text{on}\quad\partial\Omega\times(0,T].\end{cases}$
More precisely, conditions (1) and (2) above are required for a bounded function $\psi$ belonging to $C(\overline{\Omega_{T}})\,\cap L^{p}(0,T;W^{1,p}(\Omega))$. Condition (3) now reads
$\lim_{t\to 0+}\int\limits_{\Omega}\left|b(u(x,t))-b(\psi_{0}(x))\right|\hskip 0.50003pt\textup{d}x\,=\,0$
for some initial values $\psi_{0}\in C(\overline{\Omega})$. A classical example, for instance, is $\psi\equiv 1,\,\,\psi_{0}\equiv 0.$
@@ 3. Comparison principle
It will be important not to assume any existence of time derivatives. Indeed, under such an extra assumption the result is well-known. For the convenience of the reader, we present a standard proof following the notation in [LPS25a]: Consider two solutions $u_{1},u_{2}$ that have time derivatives $\partial_{t}(|u_{j}|^{p-2}u_{j})$ at least in $L^{1}(\Omega_{T})$ and take the same boundary values. Let $H_{\delta}(s)$ denote the approximation
@@
of the Heaviside function. We use the test function $\zeta=H_{\delta}\bigl(u_{2}-u_{1}\bigr)$ in both equations
@@
Next we state the comparison principle with boundary values in $C(\overline{\Omega_{T}})\cap L^{p}(0,T;W^{1,p}(\Omega))$. Observe that now no existence of time derivatives for solutions is assumed.
@@
Let $1<p<\infty.$ Let $u_{1}$ and $u_{2}$ be two weak solutions with boundary values $\psi_{1},\psi_{2}$ as in Definition 2.1. Assume that there is a constant $\varepsilon>0$ such that
(3.3) $\psi_{1}(x,t)\leq\psi_{2}(x,t)-\varepsilon\quad\text{ for all }\quad(x,t)\in\partial_{p}\Omega_{T}.$
@@
$q_{\sigma,+}(z_{1},w)-q_{\sigma,+}(z_{2},w)$ $=\,\,\int\limits_{z_{2}}^{z_{1}}\eta_{\sigma,+}^{\prime}(s-w)\,b^{\prime}(s)\,\text{d}s$
$\geq\,\,\eta_{\sigma,+}^{\prime}(z_{2}-w)\int\limits_{z_{2}}^{z_{1}}\,b^{\prime}(s)\,\text{d}s$ $=\,\,\eta_{\sigma,+}^{\prime}(z_{2}-w)\,(b(z_{1})-b(z_{2})).$
@@
By convention, we often omit the spatial variable $x$ and write $u(t)=u(x,t)$ and $\nabla u(t)=\nabla u(x,t).$ Finally, let $\phi\in C_{0}^{\infty}({\mathbb{R}})$ be a function satisfying
@@
For a parameter $\lambda>0$ and a non-negative function $\gamma\in C_{0}^{\infty}(0,T)$, we define the test function $\gamma_{\lambda}\in C_{0}^{\infty}({\mathbb{R}}^{2})$ by
@@
$|\psi_{1}(x,t)-\psi_{1}(x,s)|<\varepsilon/2\quad\text{when}\quad\left|t-s\right|<\delta$
@@
Recall that $\eta_{\sigma,+}^{\prime}(z)=0$ when $z\leq 0$. We next observe that
@@
$+$ $\gamma_{\lambda}(t,s)\eta_{\sigma,+}^{\prime\prime}(u_{1}(t)\!-\!u_{2}(s))\langle\left|\nabla u_{1}(t)\right|^{p-2}\nabla u_{1}(t),\nabla u_{1}(t)-\nabla u_{2}(s)\rangle\Big\}\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t.$
@@
As for the proof of (3.7), note that $\eta_{\sigma,-}^{\prime}(z)=0$ when $z\geq 0$. Hence, by a similar argument as in the beginning of the proof, we have that
$g(x,s):=\eta_{\sigma,-}^{\prime}(u_{2}(s)-u_{1}(t))\,\gamma_{\lambda}(t,s)$
@@
$0\,\geq\int\limits_{0}^{T}\!\!\!\int\limits_{0}^{T}\!\!\!\int\limits_{\Omega}\!$ $\Bigl\{-q_{\sigma,+}(u_{1}(t),u_{2}(s))\partial_{t}\gamma_{\lambda}(t,s)-q_{\sigma,-}(u_{2}(s),u_{1}(t))\partial_{s}\gamma_{\lambda}(t,s)\Bigr$.
@@
$\int\limits_{0}^{T}\!\!\!$ $\int\limits_{0}^{T}\!\!\!\int\limits_{\Omega}\!\left(q_{\sigma,+}(u_{1}(t),u_{2}(s))\partial_{t}\gamma_{\lambda}(t,s)+q_{\sigma,-}(u_{2}(s),u_{1}(t))\partial_{s}\gamma_{\lambda}(t,s)\right)\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t\hskip 0.50003pt\textup{d}s\,$
@@
We now prove (3.8) by letting $\sigma\to 0$. To this end, we verify a suitable convergence. First, note that when $\sigma\to 0+$
@@
$q_{\sigma,-}(z,w)\to(b(w)-b(z))^{+}\quad\text{ and }\quad\left|q_{\sigma,-}(z,w)\right|\leq(b(w)-b(z))^{+}.$
Hence, by letting $\sigma\to 0$, inequality (3.8) follows from Lebesgue’s dominated convergence theorem. ∎
Next, we let $\lambda\to 0$ in (3.8) to prove the following estimate.
@@
By a standard procedure for approximating Dirac’s delta with mollifiers, by sending $\lambda\to 0$ we arrive at
@@
See Lemma 2.9 on page 76 in [HR15] for details. This completes the proof. ∎
@@
By this and the sub-additivity of $z\mapsto z^{+}$, we get
$\int\limits_{\Omega}\left(b(u_{1}(t_{2}))-b(u_{2}(t_{2}))\right)^{+}\!\hskip 0.50003pt\textup{d}x\leq$ $\int\limits_{\Omega}\left(b(u_{1}(t_{1}))-b(\psi_{1}(t_{1}))\right)^{+}\!\hskip 0.50003pt\textup{d}x$
$+\int\limits_{\Omega}\left(b(\psi_{1}(t_{1}))-b(\psi_{2}(t_{1}))\right)^{+}\!\hskip 0.50003pt\textup{d}x$
$+\int\limits_{\Omega}\left(b(\psi_{2}(t_{1}))-b(u_{2}(t_{1}))\right)^{+}\!\hskip 0.50003pt\textup{d}x.$
@@
In some cases we can allow $\varepsilon=0$. Constant lateral boundary values are included in the next result.
Proposition 3.6 (Monotone Lateral Values).
Let $1<p<\infty.$ Suppose that the lateral boundary values are monotone, say
$\psi(x,t)\leq\psi(x,s)\quad\text{when}\quad 0<t<s<T,\,\,x\in\partial\Omega.$
Then the solution in Definition 2.1 is unique.
Proof.
We can use the same proof, if we make sure that $s\geq t$, so that $\psi(x,s)\geq\psi(x,t).$ Replace $\gamma_{\lambda}$ in (3.5) by the modified test function
$\gamma^{+}_{\lambda}(t,s)=\frac{1}{\lambda}\,\phi\Bigl(\frac{s-t}{\lambda}-1\Bigr)\,\gamma\Bigl(\frac{t+s}{2}\Bigr).$
This choice guarantees that the integrands vanish outside the interval $t<s<t+2\lambda$. Again, the function
$f(x,t)\,=\,\eta^{\prime}_{\sigma,+}\!\bigl(u_{1}(x,t)-u_{2}(x,s)\bigr)\,\gamma_{\lambda}^{+}(t,s)$
belongs to $L^{p}(0,T;W_{0}^{1,p}(\Omega)).$
@@ 4. Proof of the uniqueness
We now note that there exists one and only one such $t$-regular solution, see [LPS25a, Pages 4-6] for details. It is unique at least within their own regularity class. The next theorem guarantees that there are no other weak solutions
@@
Let $p\geq 2.$ A solution with boundary values $\psi\in C^{2}(\overline{\Omega_{T}})$ is unique.
@@
Let $u$ be an arbitrary solution with the given boundary values $\psi$. According to [LPS25a, Theorem 5], there always exists a $t$-regular solution with the boundary values $\psi\in C^{2}(\overline{\Omega_{T}})$. Moreover, it is well-known that such a $t$-regular solution is unique, see for example [LPS25a, Corollary 3], although the existence of less regular solutions is not yet excluded. Given a constant $\varepsilon>0$, there exist unique $t$-regular solutions
@@
Our convenient assumption $\psi\in C^{2}(\overline{\Omega_{T}})$ comes from the convergence argument in the proof above. To be more precise, it is based upon the finite majorant in [LPS25a, Theorem 5]
$\int\limits_{0}^{T}\!\!\!\int\limits_{\Omega}$ $\Bigl(\Big|\frac{\partial}{\partial t}\bigl(|u|^{\frac{p-2}{2}}u\bigr)\Big|^{2}+|\nabla u|^{p}+|u|^{p}\Bigr)\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t+\int\limits_{\Omega}|\nabla u(x,T)|^{p}\hskip 0.50003pt\textup{d}x$
$\leq C\int\limits_{\Omega}\left(|\psi(x,0)|^{p}+|\nabla\psi(x,0)|^{p}\right)\hskip 0.50003pt\textup{d}x$
$\hskip 50.00008pt+C\int\limits_{0}^{T}\!\!\!\int\limits_{\Omega}\Bigl(|\psi|^{p}+|\nabla\psi|^{p}+\Bigl|\frac{\partial\psi}{\partial t}\Bigr|^{p}+\Bigl|\frac{\partial\,\,}{\partial t}\nabla\psi\Bigr|^{p}\Bigr)\hskip 0.50003pt\textup{d}x\hskip 0.50003pt\textup{d}t.$
The proof also has an interesting consequence. Since we obtained $u=u_{0}$, we can read off the following result from the above estimate.
@@
$\partial_{t}(|u|^{\frac{p-2}{2}}u)\,\in L^{2}(\Omega_{T}),\qquad p\geq 2.$